# A Foundation Graph Model

### v2

@@ 4 Experiments
Num. Graphs Number of Nodes Number of Edges Diameter Avg. Clustering Task Desc.
@@
Cora 50000 59.3 $\pm$ 20.7 119$\pm$ 56.6 10.1 $\pm$ 4.68 0.322 $\pm$ 0.0845 Num. 4-Cycles
Roads 50000 59.5 $\pm$ 20.8 73.9 $\pm$ 28.2 16.4 $\pm$ 5.92 0.0559 $\pm$ 0.0426 Diameter / $|V|$
Fly Brain 50000 59.5 $\pm$ 20.8 146$\pm$ 99.6 8.8 $\pm$ 3.49 0.237 $\pm$ 0.0875 Num. 5-Cycles
@@
Community 5000 48$\pm$ 0 323$\pm$ 26.2 3$\pm$ 0.0346 0.408 $\pm$ 0.0256 Inter-Comm. Density
We begin by collating a collection of graph datasets. The datasets we use are described in aggregate measurements in Table 1. molpcba (Hu et al., 2020) A dataset of 437,929 small molecules selected from MoleculeNet (Wu et al., 2017). molesol, molfreesolv, mollipo, molclintox, molbbbp, molbace, molhiv (Hu et al., 2020) Smaller datasets sampled from MoleculeNet, with single-target classification or regression downstream tasks. Facebook (Rozemberczki et al., 2021) A single graph of page-page connections on Facebook. Originally 22,470 nodes and 171,002 edges. Downstream task set to be predicting average clustering. Twitch (Rozemberczki et al., 2020) Ego networks from the streaming platform Twitch. Ego networks have one central “hub” node, which shares an edge with all other nodes. This dataset has a downstream binary classification task. Cora (McCallum et al., 2000) A much used citation graph. 2,810 nodes and 7,981 edges. Downstream task set to be predicting the number of 4-cycles in the graph normalised by graph size. Roads (Leskovec et al., 2009) The road network of Pennsylvania. Junctions are nodes, of which there are 1,088,092, with edges roads between them, of which there are 1,541,898. Downstream task set to be predicting graph diameter normalised by graph size. Fly Brain (Winding et al., 2023) The full neural connectome of a fruit fly larvae. Each node is a neuron, of which there are 3,016. Each edge is a synapse between neurons, but as the multi-graph is dense at originally 548,000 edges, we only include an edge between neurons if there are more than two synapses between them. Downstream task set to be predicting the number of 5-cycles in the graph normalised by graph size. Trees Trees, of a randomly sampled maximum depth, and a randomly fixed probability of branching at each level. Downstream tasks is predicting depth. Random Random Erdos-Renyi (ER) graphs (Erdos and Renyi, 1960), with a number of nodes and edge probability randomly sampled. These could represent noise during training, but predicting edge probabilities from embeddings represents a useful bare-minimum downstream task during validation and testing. Community Small ER subgraphs, with a set intra-subgraph edge probability and a random inter-subgraph connection probability, with inter-subgraph connectivity as target.
As the Facebook, Cora, Roads and Fly Brain datasets consist of a single graph, we employ Exploration Sampling With Replacement (Davies and Ajmeri, 2022) to construct datasets of many smaller graphs. We use a randomly selected exploration sampler [^1] Metropolis Hasting Random Walk, Diffusion and Forest Fire samplers, from https://little-ball-of-fur.readthedocs.io/en/latest/ for each sample, which in turn samples a random number of nodes in the range $24\leq|V|\leq 96$ from the source graph. The use of a selection of exploration samplers should further ensure that a large variety of graphs and topological features are included in the datasets.

### v3

@@ 4 Experiments
Our experiments are guided by two hypotheses concerning ToP.
4.1 Datasets
We begin by collating a collection of graph datasets. Statistics are available in the Supplemental Material in Table 3. Here we provide qualitative detail as well as any downstream task information:
molpcba [14, 35]
A curated dataset of 437,929 small molecules, from which we use 250,000. Used only in pre-training.
Cora [22]
A much used citation graph. Used only in pre-training.
Fly Brain [34]
The full neural connectome of a fruit fly larvae. As the multi-graph is dense at originally 548,000 edges, we include a single edge between neurons if there are more than two synapses between them. Used only in pre-training.
Facebook [25]
A single graph of page-page connections on Facebook. Used both in pre-training and downstream transfer. On a graph level the downstream task is predicting average clustering, and we also evaluate transfer to node classification.
Twitch [26]
Ego networks from the streaming platform Twitch, used both in pre-training and downstream transfer, with a downstream binary classification task.
Roads [17]
The road network of Pennsylvania. Used both in pre-training and downstream transfer, with the task set to be predicting graph diameter normalised by graph size.
Benchmark Molecules [14]
molesol, molfreesolv, mollipo, molclintox, molbbbp, molbace, molhiv. Smaller datasets sampled from MoleculeNet, with single-target classification or regression downstream tasks. These are used only in downstream transfer.
Trees
Trees, of a randomly sampled maximum depth, and a randomly fixed probability of branching at each level, with a downstream task of predicting depth. Used only in downstream evaluation.
Random
Erdos-Renyi (ER) graphs [7], with a number of nodes and edge probability randomly sampled, and predicting edge probabilities as a downstream task. Used only in downstream evaluation.
Community
Small ER subgraphs, with a set intra-subgraph and random inter-subgraph connection probability, with inter-subgraph connectivity as target. Used only in downstream evaluation.
As the Facebook, Cora, Roads and Fly Brain datasets consist of a single graph, we employ exploration sampling to construct datasets of many smaller graphs. We use a randomly selected exploration sampler [^1] Using the littleballoffur Python library [27]. for each sample, which in turn samples a random number of nodes in the range $24\leq|V|\leq 96$ from the source graph. The use of a selection of exploration samplers should further ensure that a large variety of graphs and topological features are included in the datasets.
Figure 2: A UMAP embedding of encodings from each model, as well as an untrained model. The untrained and Chem models are noticeably more fragmented than the Social and All models. In turn the Social model is more fragmented than the All model. In the All model embedding, molecules (primarily dark blue) are all in the same region, furthest from the Twitch Ego networks (lilac), with the Cora (brown) and Facebook Page-Page data (pink) “filling the gap”.
We take the downstream task on each validation set where available and add their relevant scores (MSE and $1-$ AUROC) with a linear model as a crude monitor for the models’ representational abilities. We conduct a limited Bayesian hyperparameter sweep. The heuristic used is addition of evaluation scores on each dataset weighted by the number of samples in that dataset.
4.2 Models
We perform a limited ablation study by evaluating pre-training five models on different subsets of the amalagamated dataset: 1) GIN, an untrained GIN model of identical architecture. 2) ToP-Chem, an AD-GCL model fit only on the molecular molpcba dataset. 3) ToP-Social, an AD-GCL model fit only on non-molecular datasets (Cora, Facebook, Fly Brain, Roads, Twitch). 4) ToP-All, an AD-GCL model fit on all the training data. 5) All (CL), a model trained with the loss function from You et al. [38] using random edge dropping ($\lambda=0.2$).
Models during pre-training, unless specified, are six GIN [37] layers with single linear layers as input and output. The GIN model acts as our supervised baseline (Hypothesis 1). The ToP-Chem/ToP-Social models indicate the impact of training data diversity on downstream transfer performance, particularly on our downstream molecular datasets (Hypothesis 2). The ToP-All (CL) model shows the benefits of adversarial augmentations.
Following the results of the hyperparameter sweep, we train each model for 100 epochs with a batch size of 512, learning rates of 0.001, embedding and projection dimensions of 300 and a regularisation constant of 0.2. Figure 2 shows a UMAP embedding of encodings of the validation datasets for an untrained model and the pre-trained models. A larger version can be found in the Supplemental Material.
Pre-training was conducted on a workstation with an NVIDIA A6000, using around 16GB of video memory capacity. Pre-training the ToP-All model took a little more than eight hours. Transfer results were produced on consumer laptops, primarily with an Apple M1 processor, and on GPU use around 3GB of video memory with batch sizes of 128.
4.3 Transfer
Here we apply our pre-trained models to various downstream tasks through transfer learning. This follows stages of abstraction away from ToP pre-training: first we fine tune on datasets without features, then on datasets with features included, and finally on node and edge-level tasks.
Table 1: Scores for fine-tuning on each validation dataset. For the classification datasets we report AUROC and for regression datasets RMSE. Bold text indicates a result beats the unsupervised (GIN) baseline. Underlined text indicates a superior result, down to three significant figures. Checkmarks ✓indicate that the ToP-All model significantly out-performs the GIN baseline at $p<0.01$. ‘Best’ indicates the proportion of datasets over which a model has superior performance, and ‘Positive’ the proportion of datasets over which performance is superior to the fully-supervised baseline.
GIN (Baseline) All (CL) ToP-Chem ToP-Social ToP-All
molfreesolv 4.98 $\pm$ 0.48 4.20 $\pm$ 0.10 3.89 $\pm$ 0.21 3.83 $\pm$ 0.11 4.26 $\pm$ 0.03
molesol 1.78 $\pm$ 0.09 1.57 $\pm$ 0.06 1.44 $\pm$ 0.09 1.47 $\pm$ 0.04 1.31 $\pm$ 0.03
mollipo 1.13 $\pm$ 0.04 1.07$\pm$ 0.03 1.04 $\pm$ 0.02 0.99 $\pm$ 0.01 0.97 $\pm$ 0.01
molclintox 0.53 $\pm$ 0.04 0.46 $\pm$ 0.03 0.51 $\pm$ 0.03 0.50 $\pm$ 0.05 0.55 $\pm$ 0.11
molbbbp 0.57 $\pm$ 0.04 0.60 $\pm$ 0.03 0.54 $\pm$ 0.04 0.60 $\pm$ 0.03 0.59 $\pm$ 0.02
molbace 0.68 $\pm$ 0.05 0.76 $\pm$ 0.03 0.63 $\pm$ 0.13 0.75 $\pm$ 0.03 0.71 $\pm$ 0.02
molhiv 0.36 $\pm$ 0.06 0.50 $\pm$ 0.14 0.54 $\pm$ 0.13 0.64 $\pm$ 0.05 0.66 $\pm$ 0.04
Facebook 0.31 $\pm$ 0.07 0.24 $\pm$ 0.06 1.17 $\pm$ 0.17 0.13 $\pm$ 0.00 0.13 $\pm$ 0.00
Twitch Egos 0.72 $\pm$ 0.01 0.74 $\pm$ 0.01 0.69 $\pm$ 0.02 0.74 $\pm$ 0.01 0.79 $\pm$ 0.01
Roads 0.46 $\pm$ 0.07 0.25 $\pm$ 0.03 0.57 $\pm$ 0.07 0.10 $\pm$ 0.00 0.10 $\pm$ 0.00
Trees 0.20 $\pm$ 0.01 0.19 $\pm$ 0.02 0.24 $\pm$ 0.03 0.12 $\pm$ 0.00 0.12 $\pm$ 0.00
Community 0.74 $\pm$ 0.11 0.31 $\pm$ 0.07 0.72 $\pm$ 0.13 0.02 $\pm$ 0.00 0.02 $\pm$ 0.00
Random 0.53 $\pm$ 0.23 0.29 $\pm$ 0.10 0.59 $\pm$ 0.09 0.04 $\pm$ 0.00 0.04 $\pm$ 0.00
Best 0% 15.4% 0.0% 38.5% 69.2%
Positive 92.3% 38.5% 92.3% 100%
Features Excluded We replace the output head with a simple two-layer MLP, with a single output node. We then perform 10 fine-tuning runs on each validation dataset, with and without features, for each pre-trained model. Table 1 presents results of transfer and fine-tuning without features.
The ToP-All model consistently out-performs the other models under fine-tuning on the majority of datasets. At $p\leq 0.01$ the ToP-All model significantly out-performs the GIN baseline on 77% of datasets. When performance is not significantly better, it is never significantly worse. Notably the ToP-All model always out-performs the All (CL) model, indicating the merits of using a learnt augmentation scheme in this multi-domain setting. The standard deviation of the results on each dataset are lower than those for the non-pretrained model and the domain-specific pre-trained models. In Figure LABEL:fig:training-examples in the supplemental material we show validation loss on the several datasets during fine-tuning for the ToP-All model and an un-pre-trained GIN with the same architecture. Here the advantages of pre-training are clear: the pre-trained model has consistently lower validation loss, seems to plateau later and at a lower loss, and has a lower deviation in that loss than the baseline GIN.
Pre-training only on molecules (‘Chem’ data subsets) results at-best on-par transfer compared to the supervised GIN baseline, with negative transfer on the majority of datasets. Pre-training only on non-molecules, in comparison, results in positive transfer on the large majority of datasets and pre-training methods. At $p\leq 0.01$ the ToP-Social model significantly out-performs the ToP-Chem model on the majority of datasets. As before, where performance is not significantly better, it is never significantly worse. We employ an ablation study over the non-molecular pre-training datasets (see Supplemental Material) to ensure that this is not a single artifact result. This study corroborates the results presented in Table 1, with the Chem model consistently out-performed by non-molecular models.
We find that under pre-training with random edge dropping performance actually drops when including molecular data (‘Social’ vs ‘All’). This suggests that AD-GCL is able to differentiate between molecules and non-molecules and applies different augmentations to these domains. Nonetheless, we show that using a variety of non-molecular graphs in-general provides a more solid pre-training foundation than including molecules for both methods of augmentation.
Table 2: Scores for fine-tuning pre-trained GIN models (and a non-pre-trained baseline) on the molecular evaluation datasets with node and edge features included via projection heads. Results are reported in the same manner as in Table 1.
GIN (Baseline) All (CL) ToP-Chem ToP-Social ToP-All
molfreesolv 4.00 $\pm$ 0.43 4.00 $\pm$ 0.24 4.16 $\pm$ 0.29 3.61 $\pm$ 0.43 3.01 $\pm$ 0.59
molesol 1.72 $\pm$ 0.25 1.18 $\pm$ 0.09 1.62 $\pm$ 0.31 1.03 $\pm$ 0.17 0.90 $\pm$ 0.04
mollipo 1.09 $\pm$ 0.05 0.99 $\pm$ 0.05 1.09 $\pm$ 0.06 0.80 $\pm$ 0.02 0.79 $\pm$ 0.02
molclintox 0.65 $\pm$ 0.12 0.56 $\pm$ 0.16 0.52 $\pm$ 0.03 0.86 $\pm$ 0.02 0.84 $\pm$ 0.05
molbbbp 0.60 $\pm$ 0.03 0.63 $\pm$ 0.04 0.61 $\pm$ 0.03 0.62 $\pm$ 0.03 0.67 $\pm$ 0.02
molbace 0.75 $\pm$ 0.03 0.72 $\pm$ 0.05 0.70 $\pm$ 0.07 0.70 $\pm$ 0.05 0.72 $\pm$ 0.05
molhiv 0.32 $\pm$ 0.03 0.46 $\pm$ 0.14 0.37 $\pm$ 0.07 0.62 $\pm$ 0.11 0.57 $\pm$ 0.13
Best 14.3% 0.0% 0.0% 28.6% 57.1%
Positive 57.1% 42.9% 85.7% 85.7%
Features Included By swapping the input head of the pre-trained model, as seen in Figure 1, arbitrary node and edge features can be included during transfer. To this end we replace the input head of our model with a three layer Multi-Layer Perceptron (MLP) for both node and edge features, then perform fine-tuning runs on our molecular datasets. Table 2 presents the results of this transfer and fine-tuning. These tasks are assumed to require node and edge features to obtain good performance, and here positive transfer rates are similar to when no features are included, despite ToP’s exclusion of features during pre-training. The ToP-All and ToP-Social models again demonstrate consistent positive transfer, and on most datasets their performance improves compared to when features were excluded. The same results are presented for different GNN backbones in Table 9 in the Supplemental Material, but perform no optimisation for these models, using hyperparameters from the GIN models here for the GAT and GCN models presented. As such these results are included only for completeness.
Node & Edge Transfer By excluding the output head (again see Figure 1) we are able to transfer ToP pre-trained models onto node and edge level tasks. Here we focus on edge prediction and node classification, and though ToP pre-training does seem to offer some performance increase, the benefits are far smaller than on graph level tasks. These results can be found in the Supplemental Material in Table 10.
Further results can be found in the Supplemental Material in Section B: see Table 5 for transfer with a linear models, Tables (6,7) for GraphCL-like augmentations, Table 8 for a limited ablation over non-molecular pre-training datasets, Table 9 for ToP with varied GNN backbones, and Table 10 for node and edge classification. A comparison of ToP against domain-specific methods is in Table 11.
5 Discussion
Here we discuss our findings in terms of our hypotheses. This includes discussion on the performance of, and reasonable use cases for, models trained with ToP. Finally we detail explicitly assumptions we make and potential risks to the validity of this work.
5.1 Findings
Pre-training on multiple domains of topologies offers significant performance improvements on downstream tasks (Hypothesis 1). With features excluded we find entirely positive transfer across a range of domains and graph-level tasks using adversarially pre-trained models with ToP. With features included, on popular molecular benchmarks, there are significant performance increases on graph-level tasks compared to fully supervised models.
Further, exposure to a larger variety out-of-domain samples can benefit performance. In Table 1, assuming that the ToP-Chem model has the least varied training data, then the ToP-Social model, then the ToP-All model, performance increases along the same order (Hypothesis 2). Performance between the ToP-Social and ToP-All models is highly comparable, despite the ToP-All model having double the training samples. We expect, though we leave this as an area for future research, that a yet greater breadth of domains in the pre-training data would bring greater performance benefits.
The most intuitive assumption would be that exposing a molecular representation learner to non-molecular graphs would hinder its performance. Instead the ToP-Social and ToP-All models out-perform the ToP-Chem model almost all molecular datasets under fine-tuning. This holds true for varied pre-training regimes and model backbones (see Tables 67 and 9 in the Supplemental Material). This shows potential for future work using ToP. Pre-training on graphs could be possible through data scale and diversity in-general, instead of within a specific domain. This has clear parallels to the current trends in generalised pre-training in other fields [23, 33].
Similarly one might expect that including node and edge features during fine-tuning (or supervised training of a non-pretrained model) would likely lead to these domain-specific models out-performing all-domain models. Our results instead show that performance is both superior and more consistent from our all-domain model on featured molecular graphs, with performance increasing compared to the same model on non-featured graphs. This allows models, after using ToP, to be transferred to arbitrary tasks on graph structures, without the inherent costs of LLM-based approaches [19].
Given these findings the reasonable use-cases for models pre-trained with ToP are broad. The potential benefit is two-fold. Firstly, pre-training and fine-tuning requires less data for the downstream task to be viable. This is particularly valuable in areas of data scarcity or expense, as well as offering presumed decreases in training time for the final model. Secondly, as shown by our results in this work, our pre-trained models offer actual performance increase over simple supervised training. Our assumption is that the broad range of generic topological features learnt by the ToP model provide a more robust foundation for training than simple random initialisation, so pre-trained models avoid overfitting to non-useful features better than randomly initialised models. Guarantees on this point are left for future research.
5.2 Assumptions, Risks, Future Directions
The most significant assumption we made — that a model can learn representations that are useful on multiple domains of downstream tasks — has been justified by our results. However, in the full scope of downstream tasks for graph data, this first exploratory work is still fairly limited. Firstly we do not experiment with novel graph augmentations (or “views”), instead employing only node and edge dropping. While we’ve shown that adversarial augmentations do improve pre-training, adversarial augmentations could be developed that include some techniques present in Hassani and Khasahmadi [13], for example edge addition alongside edge dropping. Recent works have found that such augmentations, including combined additive and subtractive edge-based augmentations, can lead to more robust representation learners [9]. We anticipate that a variety of augmentations, paired with a more detailed study of how different compositions of domain data influence representations, could bolster the expressivity of these topology-pre-trained models.
Secondly our experiments are limited to a few set compositions of domains. The non-task-specific evaluation of representation models is a challenging and open research problem. The same challenge is compounded on the high-dimensional graph space, which in comparison to other data structures both lacks intuitive human understanding and has a smaller body of existing research. The usual workaround for this issue is increasing the scale and breadth of the pre-training data [28]. Future work using ToP is likely to follow this same path.
Additionally this work primarily focusses on graph-level tasks, in the same manner as the original AD-GCL work [32]. We anticipate that future work will optimise ToP, probably through targeted augmentation schemes, for node and edge related tasks specifically.
Lastly we should note that we haven’t compared ToP to other graph pre-training methods. Such methods are, with a few exceptions [19], domain specific. As such comparisons would be between our generalised models and models pre-trained on a given domain with all features included during pre-training. Such pre-training requires a large dataset of labelled samples, significantly limiting viable domains. The Supplemental Material contains a comparison of ToP pre-trained models with AD-GCL [32], GraphCL [13] and InfoGraph [36].
Liu et al. [21] recently proposed definitions for, and expected properties of, graph foundation models. The models produced with ToP satisfy their broad definition, ‘…benefit from the pre-training of broad graph data, and can be adapted to a wide range of downstream graph tasks’. The two further properties proposed by Liu et al. [21], emergence and homogenization, are less easy to demonstrate. Liu et al. themselves note that how these properties manifest is an open question, and therefore we similarly leave whether ToP models exhibit these properties as an area for future research.
6 Conclusion
We presented ToP, a method for generalised multi-domain pre-training of graph models. ToP relies on excluding node and edge features during pre-training. Features can then be re-included during fine-tuning on downstream tasks.
Findings Through contrastive learning we used ToP to present graph models capable of positive transfer across domains and tasks. These are an important step towards graph foundation models with use-cases in the manner of models such as BERT [6]. Like BERT, an early foundation language model, ToP models can be fine-tuned with positive transfer on multiple domains and tasks. On 75% of downstream tasks from multiple domains we show significant ($p\leq 0.01$) positive transfer compared to a non-pre-trained model with the same architecture, as expected from our first hypothesis. This is also true when node and edge features are re-introduced during fine-tuning. Compared to a supervised baseline, where results are not significantly better, they are never significantly worse.
Additionally we showed that ToP pre-training on a single target domain leads to at-best on-par, and on most tasks significantly worse ($p\leq 0.01$), performance compared to pre-training on multiple non-target domains. This clearly demonstrates that the variety provided by out-of-domain pre-training can hold more value than the status-quo in-domain pre-training, as in our second hypothesis.
Directions We anticipate that future work will focus on several key areas: 1) Improved or tailored graph augmentation strategies. 2) Investigation of domain weighting in dataset compositions. 3) A larger pre-training dataset spanning a wider range of domains. 4) Improved performance on node-level and edge-level tasks.
Appendix A Supplemental Material
A.1 Dataset Statistics
@@
Num. Graphs Number of Nodes Number of Edges Diameter Avg. Clustering Downstream Task
@@
Roads 50000 59.5 $\pm$ 20.8 73.9 $\pm$ 28.2 16.4 $\pm$ 5.92 0.0559 $\pm$ 0.0426 Diameter / Num. Nodes
@@
Community 5000 48$\pm$ 0 323$\pm$ 26.2 3$\pm$ 0.0346 0.408 $\pm$ 0.0256 Inter-Community Density
A.2 Exploration of Embedding Space